### v1

@@ Abstract
Supplementary information: Supplementary data are available online.
@@ 1 Introduction
High-throughput sequencing technologies have made it possible to generate large amounts of transcriptomic data, offering a wide range of materials for scientific research[17]. However, the preprocessing and analysis of these complex data presents significant challenges, particularly when it comes to building accurate and generalizable predictive models. In this article, we focus on predicting phenotypes from gene expression data. Indeed, phenotype prediction represents a major challenge in biomedical research, opening the way to a better understanding of pathological mechanisms, the identification of new biomarkers and the development of personalized therapeutic strategies.
Classical supervised machine learning methods achieve the best performance in this domain[1] and recent advances have shown that deep learning methods have great potential to outperform classical machine learning methods[20; 11], as deep neural networks can capture non-linear relationships between variables, in exchange of a large training dataset. However, these methods are highly dependent on annotated data, which can be expensive to obtain in this case. Besides, the available gene expression datasets do not contain as many examples as in other domains, like images or natural language processing where datasets are composed of millions of examples. This leads to a higher risk of overfitting for deep learning models and a limited generalization of the trained models on new datasets.
@@ 2.1 Self-supervised learning
Formally, we consider a set of labeled data $\mathcal{D}_{l}=\{x_{i},y_{i}\}_{i=1}^{|\mathcal{D}_{l}|}$ of $|\mathcal{D}_{l}|$ examples, where $x_{i}\in\mathcal{X}\subset\mathbb{R}^{N}$ denotes an example and $y_{i}\in\mathcal{Y}$ its associated label. In the context of supervised learning, we aim to learn a function $f:\mathcal{X}\rightarrow\mathcal{Y}$ by minimizing a supervised loss function (e.g. MSE or cross-entropy). With the self-supervised approach, we leverage an unlabeled dataset $\mathcal{D}_{u}=\{x_{i}\}^{|\mathcal{D}_{u}|}_{i=1}$ to learn an encoder function $e:\mathcal{X}\rightarrow\mathcal{Z}$. The space $\mathcal{Z}$ corresponds to the latent space of representations defined by a chosen pretext task. This first step is called pre-training. Generally, $|\mathcal{D}_{u}|\gg|\mathcal{D}_{l}|$ since unlabeled data is easier to collect than labeled data. Then, the learned representation space is used to solve various downstream tasks by using a small labeled dataset and attaching a classification head to the pre-trained encoder. This second step is called fine-tuning. The self-supervised training process is described in Figure 1. This approach allows a generalization, allowing the creation of robust and polyvalent models, called foundation models, which can be adapted to various downstream tasks with small annotated datasets. Self-supervised learning is mainly characterized by the strategy adopted to define the features to be predicted during the pre-training, the so-called pretext tasks. The choice of the pretext task is essential, as it impacts the representations computed by the encoder. The definition of these pretext tasks mainly relies on the exploitation of underlying domain-specific structure of the data e.g., spatial relations for images or semantic relations for the language. Therefore, the pretext task is tightly linked to the nature of the considered data. In this study, we focus on self-supervised methods for tabular data, as gene expression data are arranged in the form of high-dimension tabular data, where rows correspond to patients and columns correspond to genes. Besides, the internal structure of tabular data is the most challenging to discover as there are no explicit relationships between variables. The specificity of gene expression data is the high dimensionality ($>50$k features) and low sample size. In addition to their tabular format, some genes are highly similar and therefore are expressed similarly, which can introduce redundancy in the data.
@@ 2.2.1 SCARF
SCARF[3] is a contrastive self-supervised learning method for tabular data, described in Figure 2a. First, considering a batch of examples of the unlabeled training data, a corrupted version $\tilde{x}^{(i)}$ is generated for each example $x^{(i)}$ by sampling a fraction of the features uniformly at random and replace each of those features by a random draw from the uniform distribution over the values that feature takes on across the training dataset. Then, the pre-training process’ objective is to train an encoder $e$ that will define the representation space. The original example $x^{(i)}$ and the corrupted version $\tilde{x}^{(i)}$ pass through the encoder network $e$ to produce representations $z^{(i)}$ and $\tilde{z}^{(i)}$ that will pass through a projector $g$ to obtain projections $q^{(i)}$ and $\tilde{q}^{(i)}$ respectively. Finally, the InfoNCE contrastive loss[21] is applied to ensure that projections $q^{(i)}$ and $\tilde{q}^{(i)}$ are close, for all $i$, and projections $q^{(i)}$ and $\tilde{q}^{(j)}$ are far apart, for $i\neq j$:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}=\dfrac{1}{N}\sum_{i=1}^{N}-\log\left(\dfrac{\exp(s_{i,i}/\tau)}{\sum_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{N}\exp(s_{i,k}/\tau)}\right)$ (1)
where $s_{i,j}=q^{(i)^{\intercal}}\tilde{q}^{(j)}/(\lVert q^{(i)}\rVert_{2}.\lVert\tilde{z}^{(j)}\rVert_{2})$, for $i,j\in[N]$ and $\tau$ is the temperature hyperparameter. After the pre-training, the encoder network $e$ is kept and attached to a classification head $h$ that takes the outputs of $e$ as its inputs.
@@ 2.2.2 VIME
VIME[25] is a self-supervised generative auto-encoder-based learning method for tabular data that relies on two pretext tasks: feature reconstruction and mask reconstruction. First, the pre-training process consists in generating a binary mask vector $m=[m_{1},\cdot,m_{d}]^{\perp}\in\{0,1\}^{d}$ where $m_{j}\sim\mathcal{B}(p_{m})$, to generate a corrupted version $\tilde{x}$ for each example $x$ with the following transformation: $\tilde{x}=m\odot\overline{x}+(1-m)\odot x$, where the features of $\overline{x}$ are sampled from the empirical marginal distribution of each features.
The corrupted sample $\tilde{x}$ passes through an encoder $e$ that produces a representation $\tilde{z}$, from which a decoder $d_{f}$ predicts the values of the features that have been corrupted and a decoder $d_{m}$ predicts which features have been masked. The total loss is a combination of a feature reconstruction loss $\mathcal{L}_{f}$ and a mask reconstruction loss $\mathcal{L}_{m}$: $\mathcal{L}=\mathcal{L}_{m}+\alpha\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}_{f}$, where $\alpha$ is a hyper-parameter that adjusts the trade-off between the two losses. The pre-training process is described in Figure 2b. After the pre-training process, the encoder part $e$ is conserved and can be fine-tuned.
@@ 2.2.3 BYOL
For the pre-training process, we choose to generate a corrupted version $\tilde{x}$ of an example $x$ in the same way as VIME. Then, the online branch takes as input the original example $x$ to produce the projection $q_{\theta}(z_{\theta})$. Similarly, the target branch takes as input the corrupted examples $\tilde{x}$ to produce the representation $\tilde{z}_{\xi}$. The $l2-$distance between the representations is then minimized with the following $\mathcal{L}_{\theta,\xi}$ loss:
@@
BYOL is considered as a hybrid approach because it learns from positive pairs only. During pre-training, at each training step, the loss $\mathcal{L}_{\theta,\xi}$ is minimized with respect to $\theta$ only, as the target branch is on stop-gradient mode. On the other hand, the weights $\xi$ are updated by doing an exponential moving average of the online parameters $\theta$. More precisely, at each training step and given a target decay rate $\lambda\in[0,1]$, the following update is applied: $\xi\leftarrow\lambda\xi+(1-\lambda)\theta$. The pre-training process is described in Figure 2c. After the pre-training, only the encoder $e_{\theta}$ is conserved for the fine-tuning process.
@@ 3.1 Datasets
  • The Cancer Genome Atlas (TCGA): this is a pan-cancer RNA-Seq dataset that contains 9349 samples from 19 different types of cancer and 56902 input genes.
  • All RNA-Seq and ChIP-Seq Sample and Signature Search (ARCHS4)[15]: this is a pan-tissue samples originating from experiences from SRA (Sequence Read Archive) and GEO (Gene Expression Omnibus). It is a dataset that has not yet been widely used in the field of deep learning. It is composed of diverse samples originating from experiences that do not only focus on cancer. It therefore represents a good candidate for use as pre-training dataset. We have filtered the samples to create a dataset adapted to our experiments, composed of 45680 samples from 19 different tissue types and 67128 input genes.
@@ 3.2.1 Evaluation settings
  • ARCHS4: We divided the dataset into a pre-training set containing 50997 examples and a fine-tuning set and test set both containing 1142 examples.
@@
In each case, we take $15\%$ of both the pre-training and fine-tuning set as validation sets
@@
For SCARF, the pre-training head $g$ is a 2-layers ReLU network with hidden dimension 256. For VIME, we use the same architecture as the encoder $e$ for both the feature decoder $d_{f}$ and mask decoder $d_{m}$ with an output dimensionality of the same size of the input dimensionality. The mean square error reconstruction loss is applied for the feature decoder $d_{f}$ and the binary cross-entropy loss is applied for the mask decoder $d_{m}$. For BYOL, inspired by [7], we choose a linear layer with output size 4096 followed by batch normalization, ReLU and a final linear layer with output dimension 256 as projectors $g_{\theta}$ and $g_{\xi}$ and predictor $q_{\theta}$. The pre-training hyper-parameters are listed in Supplementary (Table S3).
@@ 3.2.2 Variation of architecture’s size
To conclude, this analysis shows that BYOL permits to build representations that use all the available dimensions, regardless of the architecture, whereas it is not the case for VIME and SCARF. However, dimensional collapse is not a sign of bad representation learning in our case, as a model that presents collapse can be used for downstream task and can still outperform the non-pre-trained version.
@@ 4 Discussion
Results show that self-supervised methods can enhance neural network performance on gene expression data, but only under certain conditions. Indeed, first the fine-tuning process has to be done with unfrozen layers, otherwise the performance of the pre-trained models does not outperform the baseline model performance. It shows that the representations induced by the three selected pre-training methods can not be used as they are and must be slightly modified through an unfrozen fine-tuning. The homogeneity of pre-training and fine-tuning datasets is crucial for the effectiveness of the pre-training step. If the datasets are homogeneous, then the pre-trained models outperform the baseline model, especially at low fine-tuning proportion $p$. This is a promising result concerning pre-training applied to gene expression data because a low fine-tuning proportion corresponds to a clinical application case for which only a few data are available ($<1000$ examples). If the distribution of pre-training and fine-tuning data are different, that may often happen in the context of batch effects, only the BYOL method is effective. SCARF and VIME fail to outperform the baseline in case of dataset heterogeneity even if we ensure that the two datasets are processed in the same way and include the same classes. BYOL is the only method that provides better performances when the pre-training and the fine-tuning set originates from different datasets, showing that the pre-training added qualitative value to the representations. The lack of improvements in this case could be due to the misalignment of the data classes. Indeed, even if the two different datasets follow the same distribution by applying the batch effect correction, those of the TCGA classes are not aligned with the distributions of the ARCHS4 classes.
In addition, the representations induced by the pre-training methods can be highly architecture-dependant. Thus, the choice of architecture is critical in order to avoid the case of counterproductive pre-training, as seen for VIME with high-layered models. However, SCARF and BYOL pre-training methods seem to be less dependent of the architecture as they only improve the baseline model by improving the performance with a few examples in fine-tuning and no counterproductive cases were encountered.
@@ 5 Conclusion
This study explores the use of self-supervised learning methods on gene expression data for phenotype prediction. We demonstrated that self-supervised learning can generate high-quality representations by exploiting the inherent structure of unlabeled data. Self-supervised learning is a cornerstone for the construction of foundation models for gene expression that may significantly improve the performance of classical supervised learning models. The three selected methods succeeded in this task, particularly when the fine-tuning is made with very few data, which is close to a real application case. Thus, these methods reduce the dependence on labeled data that is time-consuming and costly in the domain of gene expression data.
Besides, to address variability in SSL effectiveness, we provide recommendations for using these methods to observe gains compared to the baseline. Indeed, the fine-tuning should be performed on unfrozen mode. Then, if the pre-training and fine-tuning sets are homogeneous, we recommend using the SCARF method because its pre-training phase lasts less time than that of BYOL. If the sets are not homogeneous, we recommend using BYOL because it is the only method that outperforms the baseline in this case. We suggest practitioners to be cautious when using VIME as the observations indicate that its pre-training phase lasts longer and the results in fine-tuning are very unstable compared to the two other methods.

### v2

@@ Abstract
Supplementary information: Supplementary data are available at Bioinformatics online.
@@ 1 Introduction
High-throughput sequencing technologies have made it possible to generate large amounts of transcriptomic data, offering a wide range of materials for scientific research[17]. However, the preprocessing and analysis of these complex data presents significant challenges, particularly when it comes to building accurate and generalizable predictive models. In this article, we focus on predicting phenotypes from gene expression data. Phenotype prediction is key in biomedical research, aiding in understanding diseases, finding biomarkers, and developing personalized therapies.
Classical supervised machine learning methods achieve the best performance in this domain[1] and recent advances have shown that deep learning methods have great potential to outperform classical machine learning methods[20; 11], as deep neural networks can capture non-linear relationships between variables, in exchange of a large training dataset. These methods rely heavily on annotated data, which is costly to obtain. Besides, the available gene expression datasets do not contain as many examples as in other domains, like images or natural language processing where datasets are composed of millions of examples. This leads to a higher risk of overfitting for deep learning models and a limited generalization of the trained models on new datasets.
@@ 2.1 Self-supervised learning
Formally, we consider a set of labeled data $\mathcal{D}_{l}=\{x_{i},y_{i}\}_{i=1}^{|\mathcal{D}_{l}|}$ of $|\mathcal{D}_{l}|$ examples, where $x_{i}\in\mathcal{X}\subset\mathbb{R}^{N}$ denotes an example and $y_{i}\in\mathcal{Y}$ its associated label. In the context of supervised learning, we aim to learn a function $f:\mathcal{X}\rightarrow\mathcal{Y}$ by minimizing a supervised loss function (e.g. MSE or cross-entropy). With the self-supervised approach, we leverage an unlabeled dataset $\mathcal{D}_{u}=\{x_{i}\}^{|\mathcal{D}_{u}|}_{i=1}$ to learn an encoder function $e:\mathcal{X}\rightarrow\mathcal{Z}$, where $\mathcal{Z}$ is the latent space defined by a pretext task.. This first step is called pre-training. Generally, $|\mathcal{D}_{u}|\gg|\mathcal{D}_{l}|$ since unlabeled data is easier to collect than labeled data. Then, the learned representation space is used to solve various downstream tasks by using a small labeled dataset and attaching a classification head to the pre-trained encoder. This second step is called fine-tuning. The self-supervised training process is described in Figure 1. This approach allows a generalization, allowing the creation of robust and polyvalent models, called foundation models, which can be adapted to various downstream tasks with small annotated datasets. Self-supervised learning is mainly characterized by the strategy adopted to define the features to be predicted during the pre-training, the so-called pretext tasks. The choice of the pretext task is essential, as it impacts the representations computed by the encoder. The definition of these pretext tasks mainly relies on the exploitation of underlying domain-specific structure of the data e.g., spatial relations for images or semantic relations for the language. Therefore, the pretext task is tightly linked to the nature of the considered data.
In this study, we focus on self-supervised methods for tabular data, as gene expression data are arranged in the form of high-dimension tabular data, where rows correspond to patients and columns correspond to genes. Besides, the internal structure of tabular data is the most challenging to discover as there are no explicit relationships between variables. The specificity of gene expression data is the high dimensionality (>50>50k features) and low sample size. In addition to their tabular format, some genes are highly similar and therefore are expressed similarly, which can introduce redundancy in the data.
@@ 2.2.1 SCARF
SCARF[3] is a contrastive self-supervised learning method for tabular data, described in Figure 2a. First, considering a batch of examples of the unlabeled training data, a corrupted version $\tilde{x}^{(i)}$ is generated for each example $x^{(i)}$ by randomly selecting a fraction of its features. Each selected feature is replaced by a random draw from the continuous uniform distribution whose support is defined by the minimum and maximum values of that feature in the training set Then, the pre-training process’ objective is to train an encoder $e$ that will define the representation space. The original example $x^{(i)}$ and the corrupted version $\tilde{x}^{(i)}$ pass through the encoder network $e$ to produce representations $z^{(i)}$ and $\tilde{z}^{(i)}$ that will pass through a projector $g$ to obtain projections $q^{(i)}$ and $\tilde{q}^{(i)}$ respectively. Finally, the InfoNCE contrastive loss[21] is applied to ensure that projections $q^{(i)}$ and $\tilde{q}^{(i)}$ are close, for all $i$, and projections $q^{(i)}$ and $\tilde{q}^{(j)}$ are far apart, for $i\neq j$:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}=\dfrac{1}{N}\sum\limits_{i=1}^{N}-\log\left(\dfrac{\exp(s_{i,i}/\tau)}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{\begin{subarray}{c}k=1\end{subarray}}^{N}}\exp(s_{i,k}/\tau)}\right)$ (1)
where $s_{i,j}=q^{(i)^{\intercal}}\tilde{q}^{(j)}/(\lVert q^{(i)}\rVert_{2}.{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lVert\tilde{q}^{(j)}\rVert_{2}})$, for $i,j\in[N]$ and $\tau$ is the temperature hyperparameter. After the pre-training, the encoder network $e$ is kept and attached to a classification head $h$ that takes the outputs of $e$ as its inputs.
@@ 2.2.2 VIME
VIME[25] is a self-supervised generative auto-encoder-based learning method for tabular data that relies on two pretext tasks: feature reconstruction and mask reconstruction. First, the pre-training process consists of generating a binary mask vector $m=[m_{1},\cdot,m_{d}]^{\perp}\in\{0,1\}^{d}$ where $m_{j}\sim\mathcal{B}(p_{m})$, to generate a corrupted version $\tilde{x}$ for each example $x$ with the following transformation: $\tilde{x}=m\odot\overline{x}+(1-m)\odot x$, where the features of $\overline{x}$ are sampled from the empirical marginal distribution of each features.
The corrupted sample $\tilde{x}$ passes through an encoder $e$ that produces a representation $\tilde{z}$, from which a decoder $d_{f}$ predicts the values of the features that have been corrupted and a decoder $d_{m}$ predicts which features have been masked. The total loss is a combination of a feature reconstruction loss $\mathcal{L}_{f}$ and a mask reconstruction loss $\mathcal{L}_{m}$: $\mathcal{L}=\mathcal{L}_{m}+\alpha\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}_{f}$, where $\mathcal{L}_{f}$ is the mean squared error loss, $\mathcal{L}_{m}$ is the sum of the binary cross-entropy losses for each dimension of the mask vector and $\alpha$ is a hyper-parameter that adjusts the trade-off between the two losses. The pre-training process is described in Figure 2b. After the pre-training process, the encoder part $e$ is conserved and can be fine-tuned.
@@ 2.2.3 BYOL
For the pre-training process, we choose to generate a corrupted version $\tilde{x}$ of an example $x$ in the same way as VIME. Then, the online branch takes as input the original example $x$ to produce the prediction $q_{\theta}(g_{\theta}(e_{\theta}(x)))=q_{\theta}(z_{\theta})$. Similarly, the target branch takes as input the corrupted examples $\tilde{x}$ to produce the projection $g_{\xi}(e_{\xi}(\tilde{x}))=\tilde{z}_{\xi}$. The $l2$-distance between the outputs is then minimized with the following $\mathcal{L}_{\theta,\xi}$ loss:
@@
We use an overline to denote normalized quantities. BYOL is considered as a hybrid approach because it learns from positive pairs only. During pre-training, at each training step, the loss $\mathcal{L}_{\theta,\xi}$ is minimized with respect to $\theta$ only, as the target branch is on stop-gradient mode. On the other hand, the weights $\xi$ are updated by doing an exponential moving average of the online parameters $\theta$. More precisely, at each training step and given a target decay rate $\lambda\in[0,1]$, the following update is applied: $\xi\leftarrow\lambda\xi+(1-\lambda)\theta$. The pre-training process is described in Figure 2c. After the pre-training, only the encoder $e_{\theta}$ is conserved for the fine-tuning process.
@@ 3.1 Datasets
  • The Cancer Genome Atlas (TCGA): this is a pan-cancer RNA-Seq dataset that contains 9349 samples and 56902 input genes from 19 different types of cancer, with the majority class representing 12.9% of the dataset and the smallest class representing 2.8%.
  • All RNA-Seq and ChIP-Seq Sample and Signature Search (ARCHS4)[15]: this is a pan-tissue samples originating from experiences from SRA (Sequence Read Archive) and GEO (Gene Expression Omnibus). It is a dataset that has not yet been widely used in the field of deep learning. It is composed of diverse samples originating from experiences that do not only focus on cancer. It therefore represents a good candidate for use as pre-training dataset. It contains 53282 samples and 67128 input genes from 19 different tissue types, with the majority class representing 14.1% of the dataset and the smallest class representing 0.6%.
@@ 3.2.1 Evaluation settings
  • ARCHS4: We divided the dataset into a pre-training set containing 50998 examples and a fine-tuning set and test set both containing 1142 examples.
@@
In each case, we take $15\%$ of both the pre-training and fine-tuning set as validation sets. According to 3.1, a majority class classifier would achieve 12.9% of accuracy on the TCGA and ARCHS4-to-TCGA case, and 14.1% on the ARCHS4 case.
@@
For SCARF, the pre-training head $g$ is a 2-layers ReLU network with hidden dimension 256. For VIME, we use the same architecture as the encoder $e$ for both the feature decoder $d_{f}$ and mask decoder $d_{m}$ with an output dimensionality of the same size of the input dimensionality. The mean square error reconstruction loss is applied for the feature decoder $d_{f}$ and the binary cross-entropy loss is applied for the mask decoder $d_{m}$. For BYOL, inspired by [7], we choose a linear layer with output size 4096 followed by batch normalization, ReLU and a final linear layer with output dimension 256 as projectors $g_{\theta}$ and $g_{\xi}$ and predictor $q_{\theta}$. The pre-training hyper-parameters are listed in Supplementary (Table S3). We use the pre-training validation set for loss tracking during the pre-training step.
@@ 3.2.2 Variation of architecture’s size
To conclude, this analysis shows that BYOL permits to build representations that use all the available dimensions, regardless of the architecture, whereas it is not the case for VIME and SCARF. However, dimensional collapse is not a sign of bad representation learning in our case, as even collapsed models can outperform baselines on downstream task.
@@ 4 Discussion
Results show that self-supervised methods can enhance neural network performance on gene expression data, but only under certain conditions. Indeed, first the fine-tuning process has to be done with unfrozen layers, otherwise the performance of the pre-trained models does not outperform the baseline model performance. It shows that the representations induced by the three selected pre-training methods can not be used as they are and must be slightly modified through an unfrozen fine-tuning. Pre-training is most effective when datasets are homogeneous. If the datasets are homogeneous, then the pre-trained models outperform the baseline model, especially at low fine-tuning proportion $p$. This is a promising result concerning pre-training applied to gene expression data because a low fine-tuning proportion corresponds to a clinical application case for which only a few data are available (<1000<1000 examples). If pre-training and fine-tuning data differ (e.g., due to batch effects), only the BYOL method is effective. SCARF and VIME fail to outperform the baseline in case of dataset heterogeneity even if we ensure that the two datasets are processed in the same way and include the same classes. BYOL is the only method that provides better performances when the pre-training and the fine-tuning set originates from different datasets, showing that the pre-training added qualitative value to the representations. The lack of improvements in this case could be due to the misalignment of the data classes. Indeed, even if the two different datasets follow the same distribution by applying the batch effect correction, those of the TCGA classes are not aligned with the distributions of the ARCHS4 classes.
In addition, the representations induced by the pre-training methods can be highly architecture-dependant. Thus, choosing the right architecture is key to avoid counterproductive pre-training, as seen for VIME with high-layered models. However, SCARF and BYOL appear less architecture-dependent as they only improve the baseline model by improving the performance with a few examples in fine-tuning and no counterproductive cases were encountered.
@@ 5 Conclusion
This study explores the use of self-supervised learning methods on gene expression data for phenotype prediction. We showed self-supervised learning creates high-quality representations from unlabeled data. Self-supervised learning is a cornerstone for the construction of foundation models for gene expression that may significantly improve the performance of classical supervised learning models. The three selected methods succeeded in this task, particularly when the fine-tuning is made with very few data, which is close to a real application case. Thus, these methods reduce reliance on expensive, time-consuming labeled data in the domain of gene expression data.
Besides, to address variability in SSL effectiveness, we provide recommendations for using these methods to observe gains compared to the baseline. Indeed, the fine-tuning should be performed on unfrozen mode. Then, if the pre-training and fine-tuning sets are homogeneous, we recommend using the SCARF method because its pre-training phase lasts less time than that of BYOL. If the sets are not homogeneous, we recommend using BYOL because it is the only method that outperforms the baseline in this case. Practitioners should be cautious with VIME, as its pre-training is longer and the results in fine-tuning are very unstable compared to the two other methods.